kate invests money in two plans. She invests three-fifths of the money in an account at a return rate of 2%. She invests the remainder of the money in an account with a return rate of 4%. If the total interest earned in one year from investments is $84,00, how much was invested in each plan?
Answer by mananth(16949)

let total money invested be x
3x/5 ----------- 2%
2x/5-------------4%
interest
3x/5 * 0.02 + 2x/5* 0.04 = 84
0.06x/5 +0.08x/5 = 84
0.14x=5*84
0.14x=420
x= 420/0.14= 3000
2%----------- 3*3000/5 = 1800
4%------------2*3000/5 = 1200
m.ananth@hotmail.ca